OpenStudy (anonymous):

Hi guys Can I find the derivative of (5x^2 + pi*(ln(x)))/sin(square root of x) by finding the derivative of each fraction by using the quotient rule: 5x^2/sin(square root of x) + pi*(ln(x))/sin(square root of x) ? Thank you

OpenStudy (kc_kennylau):

\[\frac{5x^2+\pi\ln x}{\sin\sqrt x}\]

OpenStudy (anonymous):

Correct

OpenStudy (kc_kennylau):

Yes you can

OpenStudy (anonymous):

This is how I did it

OpenStudy (kc_kennylau):

And what did you get?

OpenStudy (anonymous):